A typical sound level for a buzzing mosquito is 35dB, and that of a vacuum cleaner is approximately 72 dB. Approximately how many buzzing mosquitos will produce a sound intensity equal to that of a vacuum cleaner?

mosquitoes
Answer #1

Sound intensity for buzzing mosquito = 35 dB = Io * 10(3.5)

Sound intensity of vacuum cleaner = 72 dB = Io * 10(7.2)

So, No. of mosquitoes needed to produce same sound level as vacuum cleaner = [Io * 10(7.2)] / [Io * 10(3.5)] = 10(7.2 - 3.5) = 10(3.7) = 5011.87